OpenStudy (anonymous):

Between what two consecutive integers does √79 lie?

OpenStudy (phi):

why don't you list out the squares of the integers 5*5= 25 6*6= 36 keep going

OpenStudy (klimenkov):

If \(0<a<b\), then \(\sqrt a<\sqrt b\). Does it help?